What is the slope and y intersept of 2x-9y=36?
Answer by stanbon(75887)

What is the slope and y intersept of 2x-9y=36?
------
Solve for "y":
9y = 2x-36
y = (2/9)x - 4
---
slope = 2/9
y-intercept = -4
==================
Cheers,
Stan H.